OpenStudy (anonymous):

can someone find D?? 10^2/2^3=8^2/D

OpenStudy (anonymous):

\[10^2/2^3=8^2/D \] becomes \[D = (2^3*8^2)/10^2\] which is \[8^3/10^2\] or \[512/100\]